# arXiv:2603.00175v1[cs.CV] 26 Feb 2026

### v1

@@ Abstract
The quadratic cost of softmax attention limits Transformer scalability in high-resolution vision. We introduce Infinite Self-Attention (InfSA), a spectral reformulation that treats each attention layer as a diffusion step on a content-adaptive token graph, accumulating multi-hop interactions through a discounted Neumann series over attention matrices. This formulation connects self-attention to classical graph centrality measures—Katz, PageRank, and eigenvector centrality—yielding interpretable and structurally grounded token weighting. We further show that this Neumann kernel coincides with the fundamental matrix of an absorbing Markov chain [46, 31], linking each token’s centrality score to its expected number of random-walk visits before absorption. Building on this, we propose Linear-InfSA, an $\mathcal{O}(N)$ variant that approximates the principal eigenvector of the implicit attention operator without forming the $N\times N$ matrix. It maintains an auxiliary state of fixed size $\mathcal{O}(d_{h})$—where $d_{h}$ is the per-head dimension, independent of the sequence length NN—is drop-in compatible with standard Vision Transformers, and supports stable forward and backward passes at $4096^{2}$ resolution and inference at $9216^{2}$ ($\sim$332k tokens). Integrated into a 4-layer ViT with 53.5M parameters and 59 GFLOPs at $224^{2}$, Linear-InfSA achieves 84.7% top-1 on ImageNet-1K, a ++3.2 pp purely architectural gain over a standard 4-layer ViT baseline (81.5%) trained with an identical recipe. On ImageNet-V2, all InfViT variants surpass every compared baseline (up to 79.8% vs. 76.8% for the best prior method), indicating robust generalization under distribution shift. Attention quality evaluations confirm semantically grounded maps: MoRF-AOC reaches 76.0% and bounding-box PR-AUC 76.1%, versus 42.6% and 56.2% respectively for softmax ViT. In scalability benchmarks on an A100 40 GB GPU, Linear-InfViT delivers 231 img/s at 0.87 J/img—a $13\times$ improvement in both throughput and energy over a standard ViT of equal depth—and is the only tested model to complete $9216^{2}$ inference without running out of memory. The linear approximation faithfully recovers the dominant eigenvector of the full quadratic operator (cosine similarity 0.985). Code and pretrained weights are available at https://github.com/giorgioroffo/infinite-self-attention.
Keywords Graph diffusion attention $\cdot$ Absorbing Markov chains $\cdot$ Linear-complexity Vision Transformers $\cdot$ Eigenvector centrality $\cdot$ Interpretable attention maps
@@ 1 Introduction
Transformer architectures underpin modern vision [13, 37, 43] and language models [60, 7, 58], yet their quadratic attention cost limits scalability in high-resolution and long-context settings [62, 10]. This has motivated numerous efficient attention mechanisms [61, 66, 11, 2]. This computational bottleneck also carries environmental cost: data-centre consumption is projected to nearly double by 2030 [29, 59], and quadratic attention dominates Transformer energy budgets [55].
@@
Despite progress, most efficient variants approximate or sparsify the attention matrix without a principled model of token interaction. Standard Transformers aggregate dependencies implicitly across stacked layers, offering limited control over multi-hop influence or interpretability. Empirical analyses further show that attention weights may highlight diffuse or semantically irrelevant regions [1, 9]. Graph-theoretic formulations offer a more structured perspective: diffusion processes quantify node influence through centrality measures such as Katz [30], PageRank [41, 3], and eigenvector centrality [5], enabling explicit multi-hop reasoning and structural interpretability—yet such principles remain underexplored within Transformer architectures. Notably, the connection between infinite-path aggregation on weighted graphs and absorbing Markov chains was established by Roffo et al. [48, 46], who showed that the fundamental matrix of a substochastic random walk yields the same Neumann kernel $N=(I-\gamma A)^{-1}$ used for centrality scoring, with entry $N_{ij}$ equal to the expected number of visits to node $j$ before absorption when starting from node $i$.
We introduce Infinite Self-Attention (InfSA), a spectral diffusion view of attention that aggregates information across layers through a truncated Neumann-series construction, akin to infinite-path kernels used in Katz/PageRank scoring [48, 46]. The same Neumann kernel admits an absorbing Markov chain interpretation [46, 31]: tokens are transient states of a random walk on the attention graph, and their centrality scores correspond to expected visit counts before diffusion terminates. We also develop Linear-InfSA, a scalable $\mathcal{O}(N)$ variant that approximates dominant attention directions via the $A^{k\to\infty}$ eigenvector limit (see Fig. 1). Integrated into Vision Transformers, InfSA delivers strong ImageNet-1K/V2 accuracy, sharper and more localized attention maps, and stable scaling to 4K–9K inputs; full derivations and extended analyses are provided in subsequent sections.
@@
The rest of the paper is organized as follows: Sec. 2 reviews related work on efficient attention, graph centrality, and state-space models; Sec. 3 derives Pure InfSA from the path-integral perspective and its absorbing Markov chain interpretation, then introduces Linear-InfSA; Sec. 4 evaluates scalability, attention quality, classification performance, ablations, eigenvector alignment, and qualitative analysis; Sec. 5 discusses scope and future directions; Sec. 6 concludes.
@@ 2 Related Work
Efficient Attention. The quadratic cost of softmax attention has driven sustained research into sub-quadratic alternatives, which can be grouped by asymptotic complexity.
$\mathcal{O}(Nd^{2})$ methods replace the $N{\times}N$ attention matrix with a product of smaller feature maps. Performer [10] approximates softmax via random Fourier features, resulting in $\mathcal{O}(Nd^{2})$ time and $\mathcal{O}(Nd)$ memory, where $d$ is the projection dimension. FLatten [21] recovers the expressiveness lost in focused linear attention through a rank-restoration mechanism whose cost also scales as $\mathcal{O}(Nd^{2})$. MLLA [22] recasts Mamba-style selective gating as linear attention, retaining an $\mathcal{O}(Nd^{2})$ bottleneck from the outer-product state update. For all three, the quadratic dependence on $d$ means that at standard head dimensions ($d_{h}{=}64$, sometimes even $d_{h}{=}128$) these methods remain noticeably slower than truly linear alternatives when $N$ is large.
$\mathcal{O}(Nd)$ methods project keys or queries to a fixed rank that does not depend on $N$. Linformer [61] projects keys and values to a fixed $k$-dimensional subspace, achieving $\mathcal{O}(Nk)$ cost; $k$ is typically set to $O(d)$. SOFT [38] applies softmax-free Gaussian kernels to achieve linear complexity without random features. Agent Attention [23] bridges softmax and linear attention through a small set of learnable agent tokens that aggregate global context, giving $\mathcal{O}(NMd)$ cost where $M\ll N$ is the fixed number of agents. Fastformer [65] replaces pairwise attention with additive pooling, which can be seen as an extreme rank-1 projection, achieving $\mathcal{O}(Nd)$ at the cost of expressiveness. While these methods scale linearly in $N$, they retain a multiplicative dependence on the head dimension $d$, which affects wall-clock performance in multi-head architectures with large $d$.
IO-optimized exact attention. FlashAttention [11, 12, 53] implements exact softmax attention through tiled, fused kernels that are aware of the GPU memory hierarchy. This eliminates the $\mathcal{O}(N^{2})$ memory wall but retains $\mathcal{O}(N^{2})$ compute: the FLOPs are unchanged, making 300k+ token sequences prohibitively slow even though they fit in memory. FlashAttention is therefore complementary to—not a replacement for—truly sub-quadratic methods.
$\mathcal{O}(N)$: Linear-InfSA. Our Linear-InfSA computes a global context vector of fixed size $\mathcal{O}(d_{h})$—independent of NN—and broadcasts it back to all tokens. The per-token cost is $\mathcal{O}(1)$ in terms of $N$, giving an overall $\mathcal{O}(N)$ complexity that is independent of $d_{h}$ in the attention computation itself. This places Linear-InfSA strictly below all of the above tiers: it avoids both the $\mathcal{O}(N^{2})$ pairwise computation of softmax attention, the $\mathcal{O}(Nd^{2})$ feature-map cost of kernel-based methods, and the $\mathcal{O}(Nd)$ projection cost of rank-reduction methods. Table 1 summarizes the complexity landscape.
Table 1: Asymptotic complexity of attention mechanisms. NN: sequence length; $d$: head dimension; kk/$M$: fixed rank / agent count. “Attn. state” refers to the per-head auxiliary memory for attention computation, excluding input/output activations which are $\mathcal{O}(Nd)$ for all methods.
Method Time Attn. state Tier
Softmax / FlashAttn [11] $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2})$ / $\mathcal{O}(N)$ Quadratic
Performer [10] $\mathcal{O}(Nd^{2})$ $\mathcal{O}(d^{2})$ Feature-map
FLatten [21] $\mathcal{O}(Nd^{2})$ $\mathcal{O}(d^{2})$ Feature-map
MLLA [22] $\mathcal{O}(Nd^{2})$ $\mathcal{O}(d^{2})$ Feature-map
Linformer [61] $\mathcal{O}(Nkd)$ $\mathcal{O}(kd)$ Projection
SOFT [38] $\mathcal{O}(Nd)$ $\mathcal{O}(d)$ Projection
Agent Attn [23] $\mathcal{O}(NMd)$ $\mathcal{O}(Md)$ Projection
Fastformer [65] $\mathcal{O}(Nd)$ $\mathcal{O}(d)$ Projection
Linear-InfSA (Ours) $\bm{\mathcal{O}(N)}$ $\bm{\mathcal{O}(d)}$ Linear
Graph Centrality, Markov Chains, and Feature Selection. The idea of interpreting pairwise interactions as a graph and deriving node importance from structural properties has a long history. Bonacich [5, 6] formalized eigenvector centrality for social networks, showing that a node’s importance is proportional to the sum of the importances of its neighbors—the dominant eigenvector of the adjacency matrix. Katz [30] proposed a related index that sums attenuated walks of all lengths, leading to the Neumann series $(I-\gamma A)^{-1}$; Page et al. [41, 3, 32] adapted this principle to the web graph as PageRank.
In the context of machine learning, Roffo et al. [48] introduced Infinite Feature Selection (Inf-FS) at ICCV 2015, recasting feature selection as a weighted-graph ranking problem. Each feature becomes a node, pairwise relations define edges, and the total importance of a feature is obtained by summing all paths of all lengths on this graph—the matrix power series $\sum_{t\geq 1}A^{t}$. To ensure convergence, the series is closed in closed form via the Neumann kernel $(I-rA)^{-1}-I$, where $r<1/\rho(A)$. This was the first work to show that infinite-path aggregation on an affinity graph can be computed in closed form and used for unsupervised feature ranking.
The theory was substantially extended in Roffo et al. [46] (IEEE TPAMI, 2021), which established that the same Neumann kernel coincides with the fundamental matrix $N=(I-Q)^{-1}$ of an absorbing Markov chain, where $Q$ is the substochastic transition matrix of the transient (feature) states. Entry $N_{ij}$ equals the expected number of visits to feature $j$ before absorption, starting from feature $i$. This Markov-chain interpretation provides a probabilistic semantics for the algebraic path-integral: feature centrality is literally the expected diffusion footprint of a random walker on the feature graph.
Roffo and Melzi [49] further connected feature ranking to eigenvector centrality in a workshop paper at ECML-PKDD 2016, observing that the dominant eigenvector of the affinity matrix assigns to each feature a centrality proportional to its recursive influence over all other features. This eigenvector perspective was formalized and experimentally validated in Roffo et al. [47], who showed that eigenvector-centrality-based ranking is robust, parameter-free (no discount $\gamma$), and competitive with the full Neumann-series approach for large feature sets.
InfSA directly inherits these foundations: we replace features with tokens, the feature-affinity matrix with the attention matrix, and the centrality ranking with token-level importance that modulates the value aggregation in each Transformer layer. Pure InfSA uses the full Neumann-series path integral; Linear-InfSA uses the eigenvector-centrality limit ($v=\lim_{k\to\infty}A^{k}x_{0}/\|A^{k}x_{0}\|_{1}$), directly inspired by the eigenvector-centrality approach of [49, 47]. This connection is not merely conceptual: the eigenvector alignment study in Sec. 4.5 shows that the Linear-InfSA weights achieve cosine similarity 0.985 with the true Perron eigenvector, empirically validating the link.
Attention as a Markov Chain: InfSA vs. TokenRank. TokenRank [14] independently interprets attention as a Markov chain but computes the stationary distribution of a closed (ergodic) chain—each token’s score is its steady-state visitation probability. InfSA instead computes the fundamental matrix $N=(I-\gamma\hat{A})^{-1}$ of an absorbing chain, encoding expected visit counts rather than steady-state probabilities. The absorbing formulation has two advantages: (i) the discount factor $\gamma$ provides explicit control over the effective diffusion depth, and (ii) the expected-visits interpretation assigns independent importance scores to each token pair, whereas a stationary distribution only provides a single scalar per token. Practically, InfSA operates as a drop-in attention replacement inside each Transformer layer, while TokenRank is a post-hoc analysis tool applied to existing attention maps.
Attention Interpretability. Attention rollout [1] propagates raw attention weights across layers; gradient-weighted attention flow [9] combines attention with gradient relevance; perturbation-based saliency [18] and MoRF/LeRF evaluation [51] assess faithfulness by progressively masking input regions. These are diagnostic tools applied post hoc to frozen attention maps. In contrast, InfSA embeds centrality directly into the attention mechanism itself, so the derived importance map is both the model’s computational routing and its explanation—unifying attention, diffusion, and attribution in a single forward pass.
@@
State-Space and Convolution Models. SSMs such as Mamba [20] and RWKV [42], along with vision adaptations MambaVision [24] and HyenaPixel [54], achieve sub-quadratic modeling via recurrence or long convolutions. These models process tokens sequentially through a hidden state, leveraging hardware-aware scan operators for efficiency. However, they do not derive token importance from graph centrality and lack the multi-hop path-integral interpretation that gives InfSA its interpretability properties. InfSA’s spectral formulation is complementary: it could in principle be layered onto SSM architectures to provide centrality-based token weighting, an avenue we leave for future work.
@@ 3 Infinite Self-Attention (InfSA)
Infinite Self-Attention (InfSA) generalizes Transformer self-attention [13] by modeling multi-hop dependencies through a spectral, path-based interpretation of token interactions, inspired by Infinite Feature Selection (Inf-FS) [48, 46]. This connects attention to classical graph diffusion and ranking—Katz, PageRank [30, 41]—and to recent spectral views of attention [50, 57, 56, 46]. We derive a layer-wise formulation compatible with standard Transformer blocks and introduce Linear-InfSA, a scalable $\mathcal{O}(N)$ variant that bypasses the $N{\times}N$ attention matrix while preserving the ViT residual structure [13].
@@ 3.1 Attention Graphs
We interpret self-attention [60, 13] as a diffusion process on a fully connected, content-adaptive graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each token is a node and each edge $(i,j)$ has weight $A_{ij}\geq 0$ derived from attention scores. Let $A\in\mathbb{R}^{N\times N}$ denote the attention matrix and $V\in\mathbb{R}^{N\times d}$ the value matrix. The self-attention update is a diffusion step:
@@
where each output token aggregates values from all others, weighted by attention [8, 64]. When $A$ is row-stochastic ($A\mathbf{1}=\mathbf{1}$), this is the random-walk operator on $\mathcal{G}$ [40]. Viewing $A$ as an affinity matrix connects attention to centrality measures—PageRank [3], Katz [30], eigenvector centrality—and to Inf-FS [48, 46], which ranks features by structural importance on weighted graphs (see Fig. 1).
@@ 3.2 Infinite Self-Attention (InfSA): Path Integrals on the Attention Graph
We extend the attention graph perspective by proposing Infinite Self-Attention (InfSA) [^1] The term “Infinite” refers to the limiting Neumann series, i.e., $\lim_{L\to\infty}\sum_{t=0}^{L}\gamma^{t}A^{t}$ and to the eigenvector limit $v=\lim_{k\to\infty}A^{k}x_{0}/\|A^{k}x_{0}\|_{1}$. Note: the actual computation is finite. We use $\gamma$ for the decay (discount) factor and reserve $\rho(\cdot)$ exclusively for the spectral radius of a matrix., a formulation inspired by the infinite path integration approach introduced by Roffo et al. [48, 46]. That approach ranks features by aggregating the weights of all paths on a feature-affinity graph, summing matrix powers $A+A^{2}+\cdots$ and closing the series via $(I{-}rA)^{-1}{-}I$. We map this construction onto self-attention: tokens replace features, the attention matrix replaces the feature-affinity matrix, and the resulting infinite-path scores become token centralities.
@@
By standard matrix algebra [26], this equals the $(i,j)$-entry of the $t$-th power of $A$:
@@
Aggregation over all path lengths. A complete assessment of token importance accounts for interactions at all depths $t=1,2,\dots$ However, the unregularized sum $\sum_{t\geq 1}A^{t}$ may diverge when $\rho(A)\geq 1$. Following [48, 46], we introduce a discount factor $\gamma\in(0,\,1/\rho(A))$ that geometrically attenuates longer paths:
@@
Closed-form via geometric matrix series. The regularized infinite-path matrix is $\check{C}=\sum_{t=1}^{\infty}(\gamma A)^{t}$. By the convergence property of geometric power series of matrices [26], if $\gamma<1/\rho(A)$ then $\rho(\gamma A)<1$ and:
@@
The proof relies on Gelfand’s formula: $\rho(\gamma A)=\rho((\gamma I)\,A)\leq\rho(\gamma I)\,\rho(A)=\gamma\,\rho(A)<1$, which guarantees $\lim_{t\to\infty}(\gamma A)^{t}=0$ and hence absolute convergence of the series [26, 19]. The matrix $\check{C}$ encodes the cumulative, discounted influence of every token on every other across all interaction depths.
When $\rho(\gamma A)<1$, the discounted path sum $\check{C}$ has a natural probabilistic reading: it coincides with the fundamental matrix of an absorbing Markov chain [46, 31]. In this view, tokens are transient states of a random walk on the attention graph, with a complementary absorption probability at each step; entry $\check{C}_{ij}+\delta_{ij}$ equals the expected number of visits to token $j$ before absorption starting from $i$. We develop this interpretation fully in Sec. 3.3.
@@
Unlike softmax, which yields row-stochastic matrices ($A\mathbf{1}{=}\mathbf{1}$) and causes oversmoothing by mixing toward a stationary distribution [34, 39], Frobenius normalization bounds the total matrix energy $\|\hat{A}\|_{F}{=}1$, providing a sufficient condition for the operator to be contractive ($\rho(\hat{A}){<}1$). This ensures the discounted series $\sum\gamma^{t}\hat{A}^{t}$ remains convergent and transforms each layer from a probability-mixing step into a centrality computation: tokens are weighted by their structural importance—akin to Katz centrality [30]—rather than by local probability mass. The absorbing Markov chain interpretation (Sec. 3.3) relies directly on this sub-stochasticity.

### v2

# Self-Attention And Beyond the Infinite: Towards Linear Transformers with Infinite Self-Attention

@@ Abstract
The quadratic cost of softmax attention limits Transformer scalability in high-resolution vision. We introduce Infinite Self-Attention (InfSA), a spectral reformulation that treats each attention layer as a diffusion step on a content-adaptive token graph, accumulating multi-hop interactions through a discounted Neumann series over attention matrices. This formulation connects self-attention to classical graph centrality measures—Katz, PageRank, and eigenvector centrality—yielding interpretable and structurally grounded token weighting. We further show that this Neumann kernel coincides with the fundamental matrix of an absorbing Markov chain [Roffo9119168, kemeny1960finite], linking each token’s centrality score to its expected number of random-walk visits before absorption. Building on this, we propose Linear-InfSA, an $\mathcal{O}(N)$ variant that approximates the principal eigenvector of the implicit attention operator without forming the $N\times N$ matrix. It maintains an auxiliary state of fixed size $\mathcal{O}(d_{h})$—where $d_{h}$ is the per-head dimension, independent of the sequence length NN—is drop-in compatible with standard Vision Transformers, and supports stable forward and backward passes at $4096^{2}$ resolution and inference at $9216^{2}$ ($\sim$332k tokens). Integrated into a 4-layer ViT with 53.5M parameters and 59 GFLOPs at $224^{2}$, Linear-InfSA achieves 84.7% top-1 on ImageNet-1K, a ++3.2 pp purely architectural gain over a standard 4-layer ViT baseline (81.5%) trained with an identical recipe. On ImageNet-V2, all InfViT variants surpass every compared baseline (up to 79.8% vs. 76.8% for the best prior method), indicating robust generalization under distribution shift. Attention quality evaluations confirm semantically grounded maps: MoRF-AOC reaches 76.0% and bounding-box PR-AUC 76.1%, versus 42.6% and 56.2% respectively for softmax ViT. In scalability benchmarks on an A100 40 GB GPU, Linear-InfViT delivers 231 img/s at 0.87 J/img—a $13\times$ improvement in both throughput and energy over a standard ViT of equal depth—and is the only tested model to complete $9216^{2}$ inference without running out of memory. The linear approximation faithfully recovers the dominant eigenvector of the full quadratic operator (cosine similarity 0.985). Code and pretrained weights will be released upon acceptance.
@@ 1 Introduction
Transformer architectures underpin modern vision [dosovitskiy2021vit, liu2022swinv2, rao2022hornet] and language models [vaswani2017attention, brown2020language, touvron2021training], yet their quadratic attention cost limits scalability in high-resolution and long-context settings [wang2020linformer, performer]. This has motivated numerous efficient attention mechanisms [linformer, xiong2021nystromformer, dao2023flashattention, beltagy2020longformer]. This computational bottleneck also carries environmental cost: data-centre consumption is projected to nearly double by 2030 [IEA2025EnergyAI, DOE2024USReport], and quadratic attention dominates Transformer energy budgets [Strubell2019EnergyNLP].
@@
Despite progress, most efficient variants approximate or sparsify the attention matrix without a principled model of token interaction. Standard Transformers aggregate dependencies implicitly across stacked layers, offering limited control over multi-hop influence or interpretability. Empirical analyses further show that attention weights may highlight diffuse or semantically irrelevant regions [abnar2020quantifying, chefer2021transformer]. Graph-theoretic formulations offer a more structured perspective: diffusion processes quantify node influence through centrality measures such as Katz [katz1953new], PageRank [page1999pagerank, bianchini2005inside], and eigenvector centrality [eigenvector_centrality], enabling explicit multi-hop reasoning and structural interpretability—yet such principles remain underexplored within Transformer architectures. Notably, the connection between infinite-path aggregation on weighted graphs and absorbing Markov chains was established by Roffo et al. [Roffo9119168], who showed that the fundamental matrix of a substochastic random walk yields the same Neumann kernel $N=(I-\gamma A)^{-1}$ used for centrality scoring, with entry $N_{ij}$ equal to the expected number of visits to node $j$ before absorption when starting from node $i$.
We introduce Infinite Self-Attention (InfSA), a spectral diffusion view of attention that aggregates information across layers through a truncated Neumann-series construction, akin to infinite-path kernels used in Katz/PageRank scoring [Roffo7410835, Roffo9119168]. The same Neumann kernel admits an absorbing Markov chain interpretation [Roffo9119168, kemeny1960finite]: tokens are transient states of a random walk on the attention graph, and their centrality scores correspond to expected visit counts before diffusion terminates. We also develop Linear-InfSA, a scalable $\mathcal{O}(N)$ variant that approximates dominant attention directions via the $A^{k\to\infty}$ eigenvector limit (see Fig. 1). Integrated into Vision Transformers, InfSA delivers strong ImageNet-1K/V2 accuracy, sharper and more localized attention maps, and stable scaling to 4K–9K inputs; derivations appear in the supplementary material.
@@
The rest of the paper is organized as follows: Sec. 2 reviews related work on efficient attention, graph centrality, and state-space models; Sec. 3 derives Pure InfSA from the path-integral perspective and its absorbing Markov chain interpretation, then introduces Linear-InfSA; Sec. 4 evaluates scalability, attention quality, and classification performance; Sec. 5 concludes.
@@ 2 Related Work
Efficient Attention. The quadratic cost of softmax attention has driven many sub-quadratic alternatives. Linformer [linformer] projects keys to fixed rank; Performer [performer] uses kernel random features; SOFT [lu2021soft] applies softmax-free kernels; FLatten [han2023flatten] restores rank in focused linear attention; Agent Attention [han2024agent] bridges softmax and linear attention via agent tokens; MLLA [han2024mlla] recasts Mamba-style gating as linear attention; and Fastformer [wu2021fastformer] replaces pairwise attention with additive pooling. FlashAttention [dao2023flashattention, dao2023flashattention2] optimizes softmax via fused kernels but retains $\mathcal{O}(N^{2})$ compute. All approximate or sparsify the attention matrix without modeling multi-hop token influence. Our Linear-InfSA computes token centrality in $\mathcal{O}(N)$ with an $\mathcal{O}(d)$ state independent of $N$, while remaining drop-in compatible with ViT blocks.
Graph Centrality and Markov Interpretations. Attention has been linked to graph reasoning [wang2018non, romero2021geometric], but few works model explicit token diffusion. We formalize attention as a content-adaptive affinity graph connecting it to Katz centrality [katz1953new], PageRank [page1999pagerank, bianchini2005inside], and Infinite Feature Selection (Inf-FS) [Roffo7410835, Roffo9119168], whose Neumann kernel coincides with the fundamental matrix of an absorbing Markov chain [kemeny1960finite]. TokenRank [erel2025attention] independently interprets attention as a Markov chain, computing the stationary distribution of a closed chain; InfSA instead computes the fundamental matrix $N{=}(I{-}\gamma\hat{A})^{-1}$ of an absorbing chain—encoding expected visit counts rather than steady-state probabilities. On interpretability, attention rollout [abnar2020quantifying], attention flow [chefer2021transformer], and perturbation saliency [fong2017interpretable] propagate relevance heuristically, while MoRF/LeRF [samek2016evaluating] evaluates faithfulness via masking. InfSA embeds centrality directly in the mechanism, unifying attention, diffusion, and attribution.
@@
State-Space and Convolution Models. SSMs such as Mamba [gu2023mamba] and RWKV [peng2023rwkv], along with vision adaptations MambaVision [hatamizadeh2025mambavision] and HyenaPixel [spravil2024hyenapixel], achieve sub-quadratic modeling via recurrence or long convolutions but do not derive token importance from graph centrality; InfSA’s spectral formulation is complementary.
@@ 3 Infinite Self-Attention (InfSA)
Infinite Self-Attention (InfSA) generalizes Transformer self-attention [dosovitskiy2021vit] by modeling multi-hop dependencies through a spectral, path-based interpretation of token interactions, inspired by Infinite Feature Selection (Inf-FS) [Roffo7410835, Roffo9119168]. This connects attention to classical graph diffusion and ranking—Katz, PageRank [katz1953new, page1999pagerank]—and to recent spectral views of attention [romero2021geometric, teo2024unveiling, kernelpca_attention, Roffo9119168]. We derive a layer-wise formulation compatible with standard Transformer blocks and introduce Linear-InfSA, a scalable $\mathcal{O}(N)$ variant that bypasses the $N{\times}N$ attention matrix while preserving the ViT residual structure [dosovitskiy2021vit].
@@ 3.1 Attention Graphs
We interpret self-attention [vaswani2017attention, dosovitskiy2021vit] as a diffusion process on a fully connected, content-adaptive graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each token is a node and each edge $(i,j)$ has weight $A_{ij}\geq 0$ derived from attention scores. Let $A\in\mathbb{R}^{N\times N}$ denote the attention matrix and $V\in\mathbb{R}^{N\times d}$ the value matrix. The self-attention update is a diffusion step:
@@
where each output token aggregates values from all others, weighted by attention [buades2005non, wang2018non]. When $A$ is row-stochastic ($A\mathbf{1}=\mathbf{1}$), this is the random-walk operator on $\mathcal{G}$ [ortega2018graph]. Viewing $A$ as an affinity matrix connects attention to centrality measures—PageRank [bianchini2005inside], Katz [katz1953new], eigenvector centrality—and to Inf-FS [Roffo7410835, Roffo9119168], which ranks features by structural importance on weighted graphs (see Fig. 1).
@@ 3.2 Infinite Self-Attention (InfSA): Path Integrals on the Attention Graph
We extend the attention graph perspective by proposing Infinite Self-Attention (InfSA) [^1] The term “Infinite” refers to the limiting Neumann series, i.e., $\lim_{L\to\infty}\sum_{t=0}^{L}\gamma^{t}A^{t}$ and to the eigenvector limit $v=\lim_{k\to\infty}A^{k}x_{0}/\|A^{k}x_{0}\|_{1}$. Note: the actual computation is finite. We use $\gamma$ for the decay (discount) factor and reserve $\rho(\cdot)$ exclusively for the spectral radius of a matrix., a formulation inspired by the infinite path integration approach introduced by Roffo et al. [Roffo7410835, Roffo9119168]. That approach ranks features by aggregating the weights of all paths on a feature-affinity graph, summing matrix powers $A+A^{2}+\cdots$ and closing the series via $(I{-}rA)^{-1}{-}I$. We map this construction onto self-attention: tokens replace features, the attention matrix replaces the feature-affinity matrix, and the resulting infinite-path scores become token centralities.
@@
By standard matrix algebra [horn2012matrix], this equals the $(i,j)$-entry of the $t$-th power of $A$:
@@
Aggregation over all path lengths. A complete assessment of token importance accounts for interactions at all depths $t=1,2,\dots$ However, the unregularized sum $\sum_{t\geq 1}A^{t}$ may diverge when $\rho(A)\geq 1$. Following [Roffo7410835, Roffo9119168], we introduce a discount factor $\gamma\in(0,\,1/\rho(A))$ that geometrically attenuates longer paths:
@@
Closed-form via geometric matrix series. The regularized infinite-path matrix is $\check{C}=\sum_{t=1}^{\infty}(\gamma A)^{t}$. By the convergence property of geometric power series of matrices [horn2012matrix], if $\gamma<1/\rho(A)$ then $\rho(\gamma A)<1$ and:
@@
The proof relies on Gelfand’s formula: $\rho(\gamma A)=\rho((\gamma I)\,A)\leq\rho(\gamma I)\,\rho(A)=\gamma\,\rho(A)<1$, which guarantees $\lim_{t\to\infty}(\gamma A)^{t}=0$ and hence absolute convergence of the series [horn2012matrix, Graham:1994]. The matrix $\check{C}$ encodes the cumulative, discounted influence of every token on every other across all interaction depths.
When $\rho(\gamma A)<1$, the discounted path sum $\check{C}$ has a natural probabilistic reading: it coincides with the fundamental matrix of an absorbing Markov chain [Roffo9119168, kemeny1960finite]. In this view, tokens are transient states of a random walk on the attention graph, with a complementary absorption probability at each step; entry $\check{C}_{ij}+\delta_{ij}$ equals the expected number of visits to token $j$ before absorption starting from $i$. We develop this interpretation fully in Sec. 3.3.
@@
Unlike softmax, which yields row-stochastic matrices ($A\mathbf{1}{=}\mathbf{1}$) and causes oversmoothing by mixing toward a stationary distribution [li2018deeper, oono2020graph], Frobenius normalization bounds the total matrix energy $\|\hat{A}\|_{F}{=}1$, providing a sufficient condition for the operator to be contractive ($\rho(\hat{A}){<}1$). This ensures the discounted series $\sum\gamma^{t}\hat{A}^{t}$ remains convergent and transforms each layer from a probability-mixing step into a centrality computation: tokens are weighted by their structural importance—akin to Katz centrality [katz1953new]—rather than by local probability mass. The absorbing Markov chain interpretation (Sec. 3.3) relies directly on this sub-stochasticity.